A rectangle is 12 cm longer than it is wide. Its perimeter is 100 cm. Find its length and width.
Answer by richard1234(7193)

Let the width be w and the length 12+w. Then,

2(w + (12+w)) = 100

w + (12+w) = 50

2w + 12 = 50 --> w = 19, so the width is 19 and the length is 31.